OpenStudy (anonymous):

When you have a number that has exactly two equal factors, each factor is called the S____ R_____ of the number. Fill in the blanks and also give an example using the proper symbol. Thanks Again!!!

OpenStudy (anonymous):

too many dashes after the r.

OpenStudy (anonymous):

I already explained this.. You have a number (lets call it N). It has two factors that are equal (lets call them n). Then \[N = n\times n = n^2\] So n = ? (in terms of N)

OpenStudy (amistre64):

square root ?

OpenStudy (anonymous):

silly rabbit

OpenStudy (amistre64):

4*4 = 16 sqrt(16) = 4

OpenStudy (anonymous):

squirrel rocket!

OpenStudy (amistre64):

summer rash?

OpenStudy (anonymous):

Its square root, @polpak and amistre64: lol?